# A Light-Cone Approach to Higher-Order Cosmological Observables

### v1

@@ 5.3 A unification of all the Geodesic Light-Cone approaches
In the gauge transformation given by Eq. (2.31), we have derived the conditions required to transform from a generic gauge fixing (later identified with the Poisson gauge) to the GLC gauge. Hence, the two systems of coordinates are related by
@@
These are indeed the expected relations in our formalism because the gauge modes to go from the PG coordinates to the GLC can be computed considering that, in general, we have
@@ 6 The Observational Angular Distance–Redshift Relation
Hereafter, we will apply the new perturbative scheme developed so far to compute the angular distance–redshift relation $d_{\text{A}}(z)$, as measured by a free-falling observer and in a perturbative and gauge-invariant approach, starting from the fully non-linear formulae (5.2) and (5.6) valid within the GLC gauge. In the evaluation of such an angular distance, as anticipated at the end of the previous section, we will go through the following steps:
@@
    First, in order to use and expand formula (5.6) in perturbation theory, we need to fix the GLC gauge at first and second order. Then we “unfix” the gauge by replacing each perturbation with each gauge-invariant counterpart provided in Eqs. (5.7) and (5.8);
@@
    Then, using the observed redshift as a time coordinate, we add to $d_{\text{A}}$ the corrections coming from expanding each quantity around the time at the source corresponding to the observed redshift, thus obtaining the fully gauge-invariant formula for the angular distance–redshift relation;
@@
    Finally, we express each light-cone perturbation in terms of the Bardeen potentials $\Phi^{(n)}$ and $\Psi^{(n)}$ and compare the results with the one already present in the literature.
In particular, we will show how the complete fixing of the residual gauge freedom (3.7) of the GLC gauge allows for the elimination of divergences at the observer position.
@@ 6.1 Perturbative expansion
$\sqrt{-g_{(4)}}$ $=a^{3}\sqrt{\bar{\gamma}}\,\bigg[1+2\nu^{(1)}+\Upsilon^{(1)}+2\nu^{(2)}+\Upsilon^{(2)}+2\Upsilon^{(1)}\nu^{(1)}+\frac{1}{2}\mathfrak{g}^{(2)}\left(\mu^{(1)},\hat{\mu}^{(1)}\right)\bigg]\,$, (6.24)
@@
$\sqrt{-g_{(4)}}$ $=a^{3}\sqrt{\bar{\gamma}}\bigg[1+2\nu^{(1)}+\frac{1}{2}N^{(1)}+2\nu^{(2)}+\frac{1}{2}N^{(2)}-\frac{1}{8}(N^{(1)})^{2}+\nu^{(1)}N^{(1)}-\frac{1}{2}U^{2}_{(1)}$
@@
As a next step, using Eqs. (6.16) and (6.18) we expand the the background and linear expression for the angular distance around the observed redshift as:
@@
$\quad-\big[\bar{\gamma}^{ab}(D_{a}\mathscr{U}^{(1)}+\tilde{D}_{a}\hat{\mathscr{U}}^{(1)})(D_{b}\mathscr{U}^{(1)}+\tilde{D}_{b}\hat{\mathscr{U}}^{(1)})\big]^{{}_{\text{o}}}_{z}-\frac{\mathscr{N}^{(1)}|^{\text{o}}_{z}\dot{\mathscr{N}}^{(1)}_{z}}{2H_{z}}\bigg\}$
@@
Now, let us compute the limit of the above quantity for $\tau\rightarrow\tau_{\text{o}}$. Since in this limit we have $r\rightarrow 0$, it follows that
@@ 6.2 Gauge fixing at the observer position
As we have already stressed, the GLC gauge at the observer position is not totally specified: this residual gauge freedom is encoded in the functions $w^{(n)}_{0}(w,\tilde{\theta}^{a})$, $\chi^{(n)}_{0}(w,\tilde{\theta}^{a})$ and $\hat{\chi}^{(n)}_{0}(w,\tilde{\theta}^{a})$ arising in the first and second-order gauge modes needed to fix the GLC gauge in perturbation theory.
@@ 6.3.1 First-order angular distance–redshift relation
$\partial_{w}X\big(\tau^{\prime},w-\eta(\tau)+\eta(\tau^{\prime})\big)$ $=\partial_{r}X(\eta^{\prime},\eta_{\text{o}}-\eta)\,\,$, (6.54)
@@ 6.3.2 Second-order angular distance–redshift relation
Thanks to the matching with the literature described in the previous paragraph, we have now validated our methodology to compute $d^{(2)}_{\text{A}}(z)$. Thus, we can systematically proceed with the evaluation of the observer terms which do not have to perfectly match with the literature, as they are due to the observational gauge fixing, correspondent to the fact that the we are computing $d_{\text{A}}(z)$ as measured by a free-falling observer.
The first is contained in Eq. (6.62), as explained at the end of the previous paragraph. The second stems from the denominator of $d^{(2)}_{\text{A}}(z)$. By Taylor expanding Eq. (6.36) and making use of
@@
Following the notation of [23], we define the second-order peculiar velocity
@@
$d^{(2)}_{\text{A}}(z)\equiv d^{(2)}_{\text{pos}}(z)+d^{(2)}_{\text{mixed}}(z)+d^{(2)}_{\text{path}}(z)+d^{(2)}_{P,\text{pos}}(z)+d^{(2)}_{P,\text{mixed}}(z)+d^{(2)}_{\text{int}}(z)\,$, (6.69)
@@
  • Finally, “int” refers to terms with peculiar velocities integrated between $\eta_{\text{in}}$ and $\eta_{\alpha}$.
@@
$d^{(2)}_{\text{pos}}(z)$ $=\bigg[1-\frac{1}{\mathcal{H}_{z}r_{z}}\bigg]\bigg[\frac{1}{2}v^{a}_{\perp z}v_{a\perp z}-\frac{1}{\mathcal{H}_{z}}(v_{||z}-v_{||\text{o}})\partial_{r}v_{||z}-(v_{||\text{o}})^{2}-v^{(2)}_{||z}\bigg$]
@@
$d^{(2)}_{\text{path,A}}(z)$ $=\left(1-\frac{1}{\mathcal{H}_{z}r_{z}}\right)\left\{-\frac{1}{r_{z}}(\Psi^{\text{A}}_{z}-\Psi^{\text{A}}_{\text{o}})\int_{\eta_{z}}^{\eta_{\text{o}}}\text{d}\eta\,\frac{\eta-\eta_{z}}{\eta_{\text{o}}-\eta}D^{2}\Psi^{\text{I}}-\Psi^{\text{A}}_{z}\Psi^{\text{A}}_{\text{o}}+(\Psi^{\text{A}}_{z})^{2}+\right$.
@@
We thoroughly compared Eqs. (6.71)-(6.75) to Eqs. (3.29)-(3.32) of [23] and found an almost perfect agreement for the purely source terms, with only few differences. One concerns the last term on the first line of (6.72) (the correspondent term of [23] has coefficient a-a which is most likely a typo, as our result agree with [20, 38]), whereas others are probably related to the way to manipulate nested integrals and deal with integration by parts. We would also like to stress that, in [54] the authors use their Eq. (A.24) to facilitate the comparison of their result (A.17) with the previous ones in the literature [20, 38]. Instead, in our case, we have shown that the corresponding relations between the coordinate perturbations in the two approaches is explicitly given in Eqs. (5.29). Indeed, at first order, our gauge modes coincide with the coordinate perturbations obtained in the first approach of [20, 38], see our Eqs. (5.21).
@@
$\quad+P_{\text{o}}\bigg\{v_{||\text{o}}\bigg(\frac{1}{2r_{z}}+\frac{\mathcal{H}_{\text{o}}}{2}-\frac{\mathcal{H}_{\text{o}}\mathcal{H}^{\prime}_{z}}{\mathcal{H}^{3}_{z}r_{z}}\bigg)+v_{||z}\bigg[\frac{1}{r_{z}}+\frac{\mathcal{H}_{\text{o}}}{\mathcal{H}_{z}r_{z}}\bigg(\frac{\mathcal{H}^{\prime}_{z}}{\mathcal{H}^{2}_{z}}-1\bigg)\bigg]$
$\quad+\bigg(1-\frac{1}{\mathcal{H}_{z}r_{z}}\bigg)\partial_{r}v_{||\text{o}}+\bigg(1-\frac{1}{\mathcal{H}_{z}r_{z}}\bigg)\bigg(1-\frac{\mathcal{H}_{\text{o}}}{\mathcal{H}_{z}}\bigg)\partial_{r}v_{||z}\bigg\}$
@@
$\quad+\bigg(-\frac{7}{2r_{z}}-\frac{1}{\mathcal{H}_{z}r^{2}_{z}}-\mathcal{H}_{\text{o}}+\frac{\mathcal{H}_{\text{o}}}{\mathcal{H}_{z}r_{z}}\bigg)\Psi^{\text{I}}_{\text{o}}-\frac{2\mathcal{H}_{\text{o}}}{r_{z}}\int_{\eta_{z}}^{\eta_{\text{o}}}\text{d}\eta\,\Psi^{\text{I}}$
@@
These last four equations represent one of the most remarkable results of this paper, because they confirm that the GLC gauge is a very convenient gauge to compute cosmological observables to higher orders in perturbation theory, having a divergences free result.
@@ 7 Summary and Conclusions
In this manuscript we have developed a second-order cosmological perturbation theory adapted to the observer past light-cone, extending previous first-order results [37]. Starting from a homogeneous and isotropic background in the Geodesic Light-Cone (GLC) coordinates, we have constructed general perturbations and analyzed their gauge transformation properties. The corresponding GLC gauge fixing conditions have been derived at second order, both in terms of light-cone perturbations and within standard perturbation theory. In particular, following [37], the standard counterpart of the GLC gauge has been dubbed Observational Synchronous Gauge (OSG).
Within the GLC gauge, the past light-cone coincides with the background one, and observables are factorized in terms of products of perturbations evaluated at the source and observer positions. Consequently, light-like cosmological observables expressed in the GLC gauge. or in the OSG one, are manifestly bi-scalar (as already stated in [30]) and depend solely on local metric perturbations. When rewritten in terms of gauge-invariant Bardeen variables, the usual integrated effects such as gravitational lensing, the ISW effect and the Shapiro time-delay naturally emerge. This provides a systematic and gauge-invariant framework to compute cosmological observables to non-linear order using the exact relations of [36, 38] and the gauge transformations connecting the GLC gauge to the Poisson Gauge (PG).
By applying this framework, we have derived the second-order expression for the angular distance–redshift relation, as an informative application and with the aim of validating our perturbative framework and method through the comparison of the results obtained to the ones of the present literature. In particular, our procedure has been based on the following steps:
  • First we identify the fully non-linear formula for the redshift and for the angular distance valid in the GLC gauge and we expand them up to second order within our perturbative framework;
  • Next, we express the angular distance in terms of the observed redshift, the observer’s past light-cone and the observed angles;
@@
The above techniques represent a new approach to compute higher-order relativistic effects to cosmological observables, validated by the fact that our results exactly reproduce [^9] A part from few terms out of hundreds, probably due to typos or differences in integration by parts and/or the way to expand terms. all known source terms from the literature based on the GLC gauge (see [20, 23, 38]). They also provide additional observer’s contributions which, to our knowledge, have not been previously identified. Indeed, our observer terms come from the fact that we have considered the angular distance–redshift relation as measured by a free-falling observer, where, working in the GLC gauge, the residual gauge freedom is exploited together with the additional requirement that the center of the coordinates coincides with the observer position. Thanks to this careful construction, non-physical divergences at the observer position have been eliminated, thus confirming the predictions and results of [30, 37].
Another important point is that, while the GLC gauge provides a simple geometrical interpretation of cosmological observables, the associated perturbation dynamics is more intricate than in standard perturbation theory. On the other hand, the OSG is fully equivalent to the GLC gauge and it is entirely formulated within the standard perturbative framework. Consequently, the dynamics and all related physical properties can be studied using the usual tools of standard perturbation theory, which may lead to an alternative solution of the long-standing problem of the dynamics in the GLC gauge. This point is left for future studies. The use of GLC guauge directly in standadrd pertubraton theory will be the subject of future studies.
In summary, we have presented a general framework to compute second-order cosmological observables, within light-cone perturbation theory (the GLC gauge). We have validated our method by recovering the known source terms in the second-order angular distance–redshift relation and identified new observer contributions, which renders the result free of divergences. This new perturbation theory can conveniently pave the way for the computation of other cosmological observables on the light-cone up to second order, as, for example, the galaxy number count [61, 62, 63, 64, 24] or the redshift drift [65, 66].

### v2

@@ 5.3 A unification of all the Geodesic Light-Cone approaches
In the gauge transformation given by Eq. (2.31), we have derived the conditions required to transform from a generic gauge fixing (later identified with the PG) to the GLC gauge. Hence, the two systems of coordinates are related by
@@
These are indeed the expected relations in our formalism because the gauge modes to go from the PG coordinates to the GLC ones can be computed considering that, in general, we have
@@ 6 The Observational Angular Distance–Redshift Relation
Hereafter, we will apply the new perturbative scheme developed so far to compute the angular distance–redshift relation $d_{\text{A}}(z)$, as measured by a free-falling observer and in a perturbative and gauge-invariant approach, starting from the fully non-linear formulae (5.2) and (5.6) valid within the GLC gauge. In the evaluation of such an angular distance, as anticipated at the end of Subsect. 5.2, we will go through the following steps:
@@
    First, we take the fully non-linear formula (5.6) valid in the GLC gauge and we expand it up to second order in the light-cone perturbations;
@@
    Second, we replace each perturbation with its gauge-invariant counterpart provided in Eqs. (5.7) and (5.8);
@@
    Then, using the observed redshift to define a new time coordinate, we add to $d_{\text{A}}$ the corrections coming from expanding each quantity around the time at the source corresponding to the observed redshift, thus obtaining the fully gauge-invariant formula for the angular distance–redshift relation;
  4. 4.
    We completely fix the observational gauge around the observer position by fixing the residual gauge freedom (3.7) of the GLC gauge. This allows for the cancellation of all the would-be divergent terms scaling as $\sim r^{-n}$ (for $n=1,2,3$) at the observer. This gauge fixing also results in further new observer terms, not present in the literature, which are needed to make the angular distance–redshift relation to be the one truly measured by a free-falling observer;
  5. 5.
    Finally, we express each light-cone perturbation in terms of the gauge-invariant Bardeen potentials $\Phi^{(n)}$ and $\Psi^{(n)}$, and compare the results with the ones already present in the literature.
Not only are the techniques described above useful to evaluate the angular distance–redshift relation, but they also constitute a general procedure to compute linear and non-linear relativistic effects of any kind of light-like cosmological observable.
@@ 6.1 Perturbative expansion
$\sqrt{-g_{(4)}}$ $=a^{4}\sqrt{\bar{\gamma}}\,\bigg[1+2\nu^{(1)}+\Upsilon^{(1)}+2\nu^{(2)}+\Upsilon^{(2)}+2\Upsilon^{(1)}\nu^{(1)}+\frac{1}{2}\mathfrak{g}^{(2)}\left(\mu^{(1)},\hat{\mu}^{(1)}\right)\bigg]\,$, (6.24)
@@
$\sqrt{-g_{(4)}}$ $=a^{4}\sqrt{\bar{\gamma}}\bigg[1+2\nu^{(1)}+\frac{1}{2}N^{(1)}+2\nu^{(2)}+\frac{1}{2}N^{(2)}-\frac{1}{8}(N^{(1)})^{2}+\nu^{(1)}N^{(1)}-\frac{1}{2}U^{2}_{(1)}$
@@
As a next step, using Eqs. (6.16) and (6.18) we expand the background and linear expression for the angular distance around the observed redshift as:
@@
$\quad-\big[\bar{\gamma}^{ab}(D_{a}\mathscr{U}^{(1)}+\tilde{D}_{a}\hat{\mathscr{U}}^{(1)})(D_{b}\mathscr{U}^{(1)}+\tilde{D}_{b}\hat{\mathscr{U}}^{(1)})\big]^{\text{o}}_{z}-\frac{\mathscr{N}^{(1)}|^{\text{o}}_{z}\dot{\mathscr{N}}^{(1)}_{z}}{2H_{z}}\bigg\}$
@@
Now, let us compute the limit of $d^{(\text{den})}_{\text{A}}(z)$ for $\tau\rightarrow\tau_{\text{o}}$. Since in this limit we have $r\rightarrow 0$, it follows that
@@ 6.2 Gauge fixing at the observer position
As we have already stressed, the GLC gauge at the observer position is not totally specified: this residual gauge freedom is encoded in the functions $w^{(n)}_{0}(w,\tilde{\theta}^{a})$, $\chi^{(n)}_{0}(w,\tilde{\theta}^{a})$ and $\hat{\chi}^{(n)}_{0}(w,\tilde{\theta}^{a})$ arising in the first- and second-order gauge modes needed to fix the GLC gauge in perturbation theory.
@@ 6.3.1 First-order angular distance–redshift relation
$\partial_{w}X\big(\tau^{\prime},w-\eta(\tau)+\eta(\tau^{\prime})\big)$ $=\partial_{r}X(\eta^{\prime},\eta_{\text{o}}-\eta)\,\,$ (6.54)
@@ 6.3.2 Second-order angular distance–redshift relation
Thanks to the matching with the literature described in the previous paragraph, we have now validated our methodology to compute $d^{(2)}_{\text{A}}(z)$. Thus, we can systematically proceed with the evaluation of the observer terms which do not have to perfectly match with the literature, as they are due to the observational gauge fixing, correspondent to the fact that we are computing $d_{\text{A}}(z)$ as measured by a free-falling observer.
The first observer contribution is contained in Eq. (6.62), as explained at the end of the previous paragraph. The second stems from the denominator of $d^{(2)}_{\text{A}}(z)$. By Taylor expanding Eq. (6.36) and making use of
@@
Following the notation of [23], we define the first-order orthogonal velocity and the second-order peculiar velocity
$v_{a\perp\alpha}$ $\equiv\partial_{a}P_{\alpha}\,$,
@@
$d^{(2)}_{\text{A}}(z)\equiv d^{(2)}_{\text{pos}}(z)+d^{(2)}_{\text{mixed}}(z)+d^{(2)}_{\text{path}}(z)+d^{(2)}_{P,\text{pos}}(z)+d^{(2)}_{P,\text{mixed}}(z)+d^{(2)}_{v,\text{int}}(z)\,$, (6.69)
@@
  • Finally, $v,\text{int}$” refers to terms with peculiar velocities integrated between $\eta_{\text{in}}$ and $\eta_{\alpha}$.
@@
$d^{(2)}_{\text{pos}}(z)$ $=\bigg[1-\frac{1}{\mathcal{H}_{z}r_{z}}\bigg]\bigg[\frac{1}{2}v^{a}_{\perp z}v_{a\perp z}-\frac{1}{\mathcal{H}_{z}}(v_{||z}-v_{||\text{o}})\partial_{r}v_{||z}-v^{(2)}_{||z}\bigg$]
@@
$d^{(2)}_{\text{path,A}}(z)$ $=\left(1-\frac{1}{\mathcal{H}_{z}r_{z}}\right)\left\{-\frac{1}{r_{z}}(\Psi^{\text{A}}_{z}-\Psi^{\text{A}}_{\text{o}})\int_{\eta_{z}}^{\eta_{\text{o}}}\text{d}\eta\,\frac{\eta-\eta_{z}}{\eta_{\text{o}}-\eta}D^{2}\Psi^{\text{I}}-\Psi^{\text{A}}_{z}\Psi^{\text{A}}_{\text{o}}+(\Psi^{\text{A}}_{z})^{2}\right$.
@@
We thoroughly compared Eqs. (6.71)-(6.75) to Eqs. (3.29)-(3.32) of [23] and found an almost perfect agreement for the purely source terms, with only few differences. One concerns the last term on the first line of (6.72) (the correspondent term of [23] has coefficient a-a which is most likely a typo, as our result agree with [20, 38]), whereas others are probably related to the way to manipulate nested integrals and deal with integration by parts. We would also like to stress that, in [54], the authors use their Eq. (A.24) to facilitate the comparison of their result (A.17) with the previous ones in the literature [20, 38]. Instead, in our case, we have shown that the corresponding relations between the coordinate perturbations in the two approaches are explicitly given in Eqs. (5.29). Indeed, at first order, our gauge modes coincide with the coordinate perturbations obtained in the first approach of [20, 38], see our Eqs. (5.21).
@@
$\quad+P_{\text{o}}\bigg\{v_{||\text{o}}\bigg(\frac{3}{2r_{z}}+\frac{\mathcal{H}_{\text{o}}}{2}-\frac{\mathcal{H}_{\text{o}}\mathcal{H}^{\prime}_{z}}{\mathcal{H}^{3}_{z}r_{z}}\bigg)+v_{||z}\bigg[\frac{1}{r_{z}}+\frac{\mathcal{H}_{\text{o}}}{\mathcal{H}_{z}r_{z}}\bigg(\frac{\mathcal{H}^{\prime}_{z}}{\mathcal{H}^{2}_{z}}-1\bigg)\bigg]$
$\quad+2\bigg(1-\frac{1}{\mathcal{H}_{z}r_{z}}\bigg)\partial_{r}v_{||\text{o}}+\bigg(1-\frac{1}{\mathcal{H}_{z}r_{z}}\bigg)\bigg(1-\frac{\mathcal{H}_{\text{o}}}{\mathcal{H}_{z}}\bigg)\partial_{r}v_{||z}\bigg\}$
@@
$\quad+\bigg(-\frac{3}{2r_{z}}-\frac{1}{\mathcal{H}_{z}r^{2}_{z}}-\mathcal{H}_{\text{o}}+\frac{\mathcal{H}_{\text{o}}}{\mathcal{H}_{z}r_{z}}\bigg)\Psi^{\text{I}}_{\text{o}}-\frac{2\mathcal{H}_{\text{o}}}{r_{z}}\int_{\eta_{z}}^{\eta_{\text{o}}}\text{d}\eta\,\Psi^{\text{I}}$
@@
These last four equations represent one of the most remarkable results of this paper, because they confirm that the GLC gauge is a very convenient gauge to compute cosmological observables to higher orders in perturbation theory, having a divergences-free result.
@@ 7 Summary and Conclusions
In this manuscript, we have developed a second-order cosmological perturbation theory adapted to the observer past light-cone, extending previous first-order results [37]. Starting from a homogeneous and isotropic background in the Geodesic Light-Cone (GLC) coordinates, we have constructed general perturbations and analyzed their gauge-transformation properties. The corresponding GLC gauge-fixing conditions have been derived at second order, both in terms of light-cone perturbations and within standard perturbation theory. In particular, following [37], the standard counterpart of the GLC gauge has been dubbed Observational Synchronous Gauge (OSG).
Within the GLC gauge, the past light-cone coincides with the background one, and observables are factorized in terms of products of perturbations evaluated at the source and observer positions. Consequently, light-like cosmological observables expressed in the GLC gauge, or in the OSG one, are manifestly bi-scalar (as already stated in [30]) and depend solely on local metric perturbations. When rewritten in terms of gauge-invariant Bardeen variables, the usual integrated effects such as gravitational lensing, the ISW effect and the Shapiro time-delay naturally emerge. This provides a systematic and gauge-invariant framework to compute cosmological observables to non-linear orders using the exact relations of [36, 38] and the gauge transformations connecting the GLC gauge to the Poisson Gauge (PG).
By applying this framework, we have derived the second-order expression for the angular distance–redshift relation, as an informative application and with the aim of validating our perturbative framework and method through the comparison of our results obtained with the ones of the present literature. In particular, our procedure has been based on the following steps:
  • First we identify the fully non-linear formulae for the redshift and for the angular distance valid in the GLC gauge and we expand them up to second order within our perturbative framework;
  • Next, we express the angular distance in terms of the observed redshift, the observer past light-cone and the observed angles;
@@
The above techniques represent a new approach to compute higher-order relativistic effects to cosmological observables, validated by the fact that our results exactly reproduce [^9] A part from few terms out of hundreds, probably due to typos or differences in integration by parts and/or the way to expand terms. all known source terms from the literature based on the GLC gauge (see [20, 23, 38]). They also provide additional observer contributions which, to our knowledge, have not been previously identified. Indeed, our observer terms come from the fact that we have considered the angular distance–redshift relation as measured by a free-falling observer, where, working in the GLC gauge, the residual gauge freedom is exploited together with the additional requirement that the center of the coordinates coincides with the observer position. Thanks to this careful construction, non-physical divergences at the observer position have been eliminated, thus confirming the predictions and results of [30, 37].
Another important point is that, while the GLC gauge provides a simple geometrical interpretation of cosmological observables, the associated perturbation dynamics is more intricate than in standard perturbation theory. On the other hand, the OSG is fully equivalent to the GLC gauge and it is entirely formulated within the standard perturbative framework. Consequently, the dynamics and all related physical properties can be studied using the usual tools of standard perturbation theory, which may lead to an alternative solution of the long-standing problem of the dynamics in the GLC gauge. The use of GLC gauge directly in standard perturbation theory will be the subject of future studies.
In summary, we have presented a general framework to compute second-order cosmological observables, within the light-cone perturbation theory (the GLC gauge). We have validated our method by recovering the known source terms in the second-order angular distance–redshift relation and identified new observer contributions, which renders the result free of divergences. This new perturbation theory can conveniently pave the way for the computation of other cosmological observables on the light-cone up to second order, as, for example, the galaxy number count [61, 62, 63, 64, 24] or the redshift drift [65, 66].